Weekly Inspection- 26th April


Weather – more of the same – wet and cold. Temperatures not getting above 14°C. Forecast to warm up next week.

Forage: hawthorn, OSR, Apple, Green alkanet and lilac just starting.

Colonies starting to expand- still very early in the season. Will need to decide on where to put the new colony being created by Blue using the modified Snelgrove method which seems to be progressing well.

Hive Record
CommentNext Inspection
AquaYY10 frames of BIAS spread across 2 boxesPlenty in the boxesUnite gone well with plenty of broodAdd 4 frames if staying with double brood boxes
BlueNY4 in queen right bottom box, 5 capped in top box with 12 QCs - took down to 1Plenty in the brood boxDidn’t see queen or eggs - need to check next weekLook for mated queen in 10 days
CyanNY5 BIAS - Not sure if they are expanding Super being worked - nothing stored yet.Not sure whether queen is present. Seem to be fewer beesCheck for queen
DenimNN 4 frames capped brood, 4 capped QCs - took down to 1Plenty in the brood boxLooks to be OK with raising new queenCheck in 10 days for mated queen
EmeraldYY3 frames of BIAS - was 2 frames before when Nuc createdBringing in nectarNot many beesCheck on expansion - strengthen if necessary
FlaxYY4 frames of BIAS, continuing to expandPlenty of stores Settled into new site very wellReplace 3 old 14 x 12 frames
HoneydewNY5 frames of BIAS, - plenty of eggs so expanded well in last 2 weeksNot much in the brood box, but bringing in nectarAdded super as weather set to improveCheck on continued expansion and find queen